If Y1, Y2 are identically independently distributed as normal distribution with mean u


1
Expert's answer
2021-06-21T12:51:21-0400

Let "Y_1" and "Y_2" are independent and identically distributed normal random variable with mean "\mu" and variance "\sigma^2."

It is known that a linear combination of independent normal random variables is also normally distributed.

If "Y_1" and "Y_2" are independent and identically distributed normal random variable with mean "\mu" and variance "\sigma^2," then "U=\dfrac{1}{2}(Y_1-3Y_2)" is also normally distributed. So, the mean and variance of "U"are,


"\mu_U=E[U]=E[\dfrac{1}{2}(Y_1-3Y_2)]=\dfrac{1}{2}E[Y_1-3Y_2]"

"=\dfrac{1}{2}(E[Y_1]-3E[Y_2])=\dfrac{1}{2}(\mu-3\mu)=-\mu"

"\sigma^2_U=V[U]=V[\dfrac{1}{2}(Y_1-3Y_2)]"

"=\dfrac{1}{4}(V[Y_1]+(-3)^2V[Y_2])"

"=\dfrac{1}{4}(\sigma^2+9\sigma^2)=\dfrac{5}{2}\sigma^2"

Therefore, "U" follows a normal distribution with mean "-\mu" and variance "\dfrac{5}{2}\sigma^2"


"U\sim N\big(-\mu, \dfrac{5}{2}\sigma^2\big)"